If (5x−2)^2=ax^2−bx+c, what is the value of a+c^2?
Answer by stanbon(75887)

If (5x−2)^2=ax^2−bx+c, what is the value of a+c^2?
---------
25x^2-20x+4 = ax^2-bx+c
------
a = 25 ; b = 20 ; c = 4
----
Ans: a+c^2 = 25+4^2 = 25+16 = 41
---------------
Cheers,
Stan H
----------------